# ProjectTest: A Project-level LLM Unit Test Generation Benchmark and Impact of Error Fixing Mechanisms

### v2

@@ Abstract
Unit test generation has become a promising and important use case of LLMs. However, existing evaluation benchmarks for assessing LLM unit test generation capabilities focus on function- or class-level code rather than more practical and challenging project-level codebases. To address such limitation, we propose ProjectTest, a project-level benchmark for unit test generation covering Python, Java, and JavaScript. ProjectTest features 20 moderate-sized and high-quality projects per language. We evaluate nine frontier LLMs on ProjectTest and the results show that all frontier LLMs tested exhibit moderate performance on ProjectTest on Python and Java, highlighting the difficulty of ProjectTest. We also conduct a thorough error analysis, which shows that even frontier LLMs, such as Claude-3.5-Sonnet, have significant simple errors, including compilation and cascade errors. Motivated by this observation, we further evaluate all frontier LLMs under manual error-fixing and self-error-fixing scenarios to assess their potential when equipped with error-fixing mechanisms.
@@ 1 Introduction
Given the rapid adoption of LLMs on unit testing, the evaluation of LLM unit test generation capabilities appears to be lagging behind. Previous unit test generation evaluation benchmarks primarily focus on function-level, class-level, or file-level Chen et al. (2021); Du et al. (2023); Wang et al. (2024); Jain et al. (2024a) codes. However, project-level codes are more representative of real-world scenarios and practical needs. The complex dependency relationships between different files in project-level codebases make unit test generation more challenging. The only existing benchmark that has briefly explored project-level unit test generation is DevBench Li et al. (2024). However, due to its broad focus, the number of projects included for unit test generation is low for each language (e.g., 5 for Java and 5 combined for C and C++), with varying quality. Half of its projects for unit test generation evaluation are difficulty to track, and most of the identifiable projects have fewer than 250 Stars and fewer than 50 Forks. DevBench also didn’t provide a thorough analysis of error types, potentials, or self-fixing capabilities of frontier LLMs’ project-level unit test generation.
@@
Error analyses from above also show that even frontier LLMs, like Claude-3.5-Sonnet, have significant compilation and cascade errors. Although these errors appear to be preliminary and may be relatively easy to fix, they prevent us from observing more advanced aspects of LLM performance on unit test generation, such as correctness and coverage. To address this, we first manually fix LLM’s compilation and cascade errors and then re-evaluate the fixed unit tests. This allows us to measure not only the models’ raw performance but also their potential for improvement when combined with simple error-fixing mechanisms. By incorporating error-fixing, we uncover critical insights into the effort required to refine generated tests and better understand the various types of errors that occur in unit tests generated by different LLMs. We observe that the model rankings change significantly after the manual fix, showing the significant differences in different LLMs’ error distribution and their potentials after error-fixing. Inspired by such findings from manual fixes, we also explore using LLMs for self-fixing their errors in generating project-level unit tests. The results show that while LLMs can correct some errors in their generated unit tests, their self-fixing abilities still lag behind the quality and reliability of human fixes.
We summarize the contribution of this paper as follows: we introduce the first project-level evaluation benchmark for unit test generation and conduct an extensive evaluation of nine frontier LLMs. Additionally, we conduct thorough error analyses by manually fixing compilation and cascade errors and provide critical insights. Inspired by the error analysis, we are the first to assess LLMs’ self-fixing capability on unit test generation.
@@ 2.3 Unit Test Generation Benchmark
Current benchmarks for LLM-based unit test generation mainly focus on function-level Wang et al. (2024), class-level Du et al. (2023), or file-level code Jain et al. (2024a). Project-level software testing benchmarks, on the other hand, often target tasks other than unit test generation. For instance, R2E-Eval1 Jain et al. (2024b) is designed for the generation of equivalent test harnesses, SWT-Bench Mündler et al. (2024) focuses on fixing specific bugs rather than entire projects, and DevBench Li et al. (2024) centers on software development tasks. While DevBench skims the surface of project-level unit testing, its data size and quality are quite mixed, particularly for C/C# and Java, with only five projects, respectively. Additionally, given its broad focus, DevBench doesn’t include comprehensive evaluation and error analysis on project-level unit test generation.
@@ 3 Methodology
We first introduce the dataset collection and preprocessing of creating ProjectTest (§3.1). Then, we introduce evaluation metrics (§3.2) and provide an overview of the Unit Test Generation pipeline (§3.3) that we use to evaluate LLMs on ProjectTest, including three unit test generation scenarios.
@@ 3.1 Benchmark Dataset
Dataset Collection
Our dataset is built from carefully selected project-level repositories on GitHub, which is a popular platform for hosting and collaborating on software development projects. We focus on three widely used programming languages: Python, Java, and JavaScript. We establish our selection criteria based on three key factors: 1) a reasonable size, 2) inter-file dependencies, and 3) a reliable source. Thus, we collect reliable and self-contained projects consisting of 2 to 15 files with fewer than 1600 lines of code (LOC). We limit our selection to repositories with publicly available licenses, such as the MIT license, ensuring the legality and openness of the code. To maintain the quality and reliability of the dataset, we choose projects with a high number of stars and forks, which signals community approval and widespread usage. For projects that fit all requirements above but are too big for current frontier LLMs to handle, we also extracted smaller projects from these large codebases. These smaller projects were carefully adjusted to be self-contained without relying on the original larger projects. After applying these criteria, we constructed 20 representative projects per programming language. The summary of dataset statistics is presented in Table 1, and detailed information on the dataset sources and statistics for each project can be found in Appendix A.
Pre-processing
Dataset pre-processing involves several key steps to ensure that the projects are well-structured and suitable for testing. First, we double-check whether the selected projects have syntax errors, even though they are sourced from reliable codebases. Second, for projects extracted from a larger codebase, we modify them to be self-contained by reorganizing files, adjusting domain naming conventions, and/or modifying import paths to remove dependencies on external modules. Next, to enhance the accuracy of line coverage measurements, we consolidate statements that are split across multiple lines into a single line, ensuring that the metrics are more valid. Additionally, we maintain the integrity of the original code style as much as possible, preserving the diverse coding practices across different projects. This approach allows us to test how LLMs handle various code styles in a realistic environment.
@@ 3.2 Evaluation Metrics
We focus on three key aspects when evaluating the generated unit tests: compilation rate, correctness rate, and coverage rate. Compilation rate (ComR) measures the percentage of projects in which the generated test suites compile successfully, indicating how often LLMs produce unit test suites that can be executed without compilation errors. The compilation rate for all projects in $X$ is defined as $ComR=\frac{|X^{com}|}{|X|}$, where $X$ is the project set and $X^{com}\subset X$ denotes the subset of projects whose test suites compile successfully. Correctness rate (CR) calculates the percentage of unit tests that are correct out of all generated unit tests for each project, providing insight into the accuracy of the test generation process. On average, more than 95% of vanilla-generated unit tests compare expected and actual values, reinforcing the validity of CR as an evaluation metric. Detailed statistics see Appendex C. The correctness rate for the project $x$ is defined as $CR_{x}=\frac{|T_{x}^{cor}|}{|T_{x}|}$, where $T_{x}$ is the generated test suite and $T_{x}^{cor}\subset T_{x}$ denotes the correct unit test set for the project $x$. Coverage rate analyzes both line and branch coverage to understand how well the generated unit tests explore the code’s functionality. The coverage rate for the project $x$ is defined as $CR_{x}=\frac{covered(x)}{total(x)}$, where $covered(x)$ denotes the number of covered lines/branches in project $x$ and $total(x)$ the total number of lines/branches in project $x\in X$.
@@ 3.3 Unit Test Generation
Inspired by our observation from manual fixing that different LLMs exhibit significantly different potentials after manual fixing, we are curious about how LLMs would perform in self-fixing on our benchmark. We explore LLMs’ self-fixing abilities by incorporating conversation history and error messages as shown in Figure 6. We provide LLMs with the conversation history (including the system prompt, the user prompt for unit test generation requests, and LLM vanilla response), error messages obtained from the testing framework, and the user prompt for error fixing requests. When the open-source LLM’s input length is limited, we prioritize the information in the following order: system prompt, LLM’s initial response, error messages, error-fixing requests, and unit test generation requests. Less important information is truncated as needed. Additionally, we reserve at least 2,000 tokens for the open-source LLM’s self-fixing outputs. LLM self-fixing scenario helps us understand LLMs’ error-fixing ability and their potential to generate better unit tests when incorporating the self-fixing process. Note that during self-fixing, we do not constrain the target error types to just compilation or cascade errors.
@@ 5.1 Main Results
Table 2 shows the evaluation results of the vanilla unit tests. First, we observe that different LLMs have varying language-level expertise. For example, Claude-3.5-Sonnet performs the best in Java but falls behind GPT-o1 on JavaScript. Second, we can see from the results that LLMs having different metric-level expertise as well, validating the effectiveness of different evaluation metrics. For example, in Python, Claude-3.5-Sonnet performs the best on CR and ComR while falling behind GPT-o1 on LC and BC.
Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Many models fail to generate valid Java code, leading to low compilation rates and execution coverage. Among all the evaluated models, GPT-o1 performs the best in general, especially in JavaScript. CodeLlama and CodeGemma have the worst general performance. We also observe that some models tend to generate more unit tests. However, generating more unit tests does do not necessarily lead to better coverage rates. For example, Gemini-1.5-Pro tends to generate the most unit tests but does not obtain the best coverage rate. Additionally, we observe that sometimes the open-source model can even outperform some closed-source models. For example, DeepSeek-Coder works better than GPT-3.5-Turbo on Python and JavaScript. Finally, we confirmed from such results that dependencies exist in metrics. On Java, models like CodeQwen1.5, CodeLlama, and CodeGemma fail to generate compilable unit tests, resulting in the lowest correctness rates and coverage rates.
@@ 5.4 Unique Contribution of Unit Tests
We also explore the unique contribution of the generated unit tests on Python. The unique contribution is defined as the total portion of coverage contributed by each generated unit test that does not overlap with the coverage of other unit tests. This is important for several reasons. First, some LLMs generate more unit tests than others, making it insufficient to rely solely on coverage rate as a metric; the unique contribution of each test should also be considered. Second, it’s crucial for LLMs to generate fewer unit tests while still achieving a high coverage rate, as running a large number of tests can sometimes be resource- or time-intensive.

### v3

@@ Abstract
Unit test generation has become a promising and important use case of LLMs. However, existing evaluation benchmarks for assessing LLM unit test generation capabilities focus on function- or class-level code rather than more practical and challenging project-level codebases. To address such limitation, we propose ProjectTest, a project-level benchmark for unit test generation covering Python, Java, and JavaScript. ProjectTest features 20 moderate-sized and high-quality projects per language. We evaluate nine frontier LLMs on ProjectTest and the results show that all frontier LLMs tested exhibit moderate performance on ProjectTest on Python and Java, highlighting the difficulty of ProjectTest. We also conduct a thorough error analysis, which shows that even frontier LLMs, such as Claude-3.5-Sonnet, have significant basic yet critical errors, including compilation and cascade errors. Motivated by this observation, we further evaluate all frontier LLMs under manual error-fixing and self-error-fixing scenarios to assess their potential when equipped with error-fixing mechanisms. Our code and dataset is available at ProjectTest.
@@ 1 Introduction
Given the rapid adoption of LLMs on unit testing, the evaluation of LLM unit test generation capabilities appears to be lagging behind. Previous unit test generation evaluation benchmarks primarily focus on function-level, class-level, or file-level Chen et al. (2021); Du et al. (2023); Wang et al. (2024); Jain et al. (2024a) codes. However, project-level codes are more representative of real-world scenarios and practical needs. The complex dependency relationships between different files in project-level codebases make unit test generation more challenging. The only existing benchmark that has briefly explored project-level unit test generation is DevBench Li et al. (2024). However, due to its broad focus, the number of projects included for unit test generation is low for each language (e.g., 5 for Java and 5 combined for C and C++), with varying quality. Half of its projects for unit test generation evaluation are difficult to track, and most of the identifiable projects have fewer than 250 Stars and fewer than 50 Forks. DevBench also does not provide a thorough analysis of error types, potentials, or self-fixing capabilities of frontier LLMs’ project-level unit test generation.
@@
Error analyses from above also show that even frontier LLMs, like Claude-3.5-Sonnet, have significant compilation and cascade errors. Although these errors appear to be preliminary and may be relatively easy to fix, they prevent us from observing more advanced aspects of LLM performance on unit test generation, such as correctness and coverage. To address this, we first manually fix LLM’s compilation and cascade errors and then re-evaluate the fixed unit tests. This allows us to measure not only the models’ raw performance but also their potential for improvement when combined with error-fixing mechanisms. By incorporating error-fixing, we uncover critical insights into the effort required to refine generated tests and better understand the various types of errors that occur in unit tests generated by different LLMs. We observe that the model rankings change significantly after the manual fix, showing the significant differences in different LLMs’ error distribution and their potentials after error-fixing. Inspired by such findings from manual fixes, we also explore using LLMs for self-fixing their errors in generating project-level unit tests. The results show that while LLMs can correct some errors in their generated unit tests, their self-fixing abilities still lag behind the quality and reliability of human fixes.
We summarize our contributions as follows: we introduce the first project-level evaluation benchmark for unit test generation and conduct an extensive evaluation of nine frontier LLMs. Additionally, we conduct thorough error analyses by manually fixing compilation and cascade errors and provide critical insights. Inspired by the error analysis, we are the first to assess LLMs’ self-fixing capability on unit test generation.
@@ 2.3 Unit Test Generation Benchmark
Current benchmarks for LLM-based unit test generation mainly focus on function-level Wang et al. (2024), class-level Du et al. (2023), or file-level code Jain et al. (2024a). Project-level software testing benchmarks, on the other hand, often target tasks other than unit test generation. For instance, R2E-Eval1 Jain et al. (2024b) is designed for the generation of equivalent test harnesses, SWT-Bench Mündler et al. (2024) focuses on fixing specific bugs rather than entire projects, and DevBench Li et al. (2024) centers on software development tasks. While DevBench touches on project-level unit testing, its dataset is limited in quantity and varies in quality, especially for C/C# and Java, with only five projects each. Moreover, due to its broad focus, DevBench lacks a comprehensive evaluation and error analysis of LLM project-level unit test generation.
@@ 3 Methodology
We first introduce the dataset collection and preprocessing of creating ProjectTest (§3.1). Then, we introduce evaluation metrics (§3.2) and the Unit Test Generation pipeline (§3.3) that we use to evaluate LLMs on ProjectTest, including three unit test generation scenarios.
@@ 3.1 Benchmark Dataset
Dataset Collection. Our dataset is built from carefully selected project-level repositories on GitHub, which is a popular platform for hosting and collaborating on software development projects. We focus on three widely used programming languages: Python, Java, and JavaScript. We establish our selection criteria based on three key factors: 1) a reasonable size, 2) inter-file dependencies, and 3) a reliable source. Thus, we collect reliable and self-contained projects consisting of 2 to 15 files with fewer than 1600 lines of code (LOC). We limit our selection to repositories with publicly available licenses, such as the MIT license, ensuring the legality and openness of the code. To maintain the quality and reliability of the dataset, we choose projects with a high number of stars and forks, which signals community approval and widespread usage. For projects that fit all the requirements above but are too big for current frontier LLMs to handle, we also extracted smaller projects from these large codebases. These smaller projects were carefully adjusted to be self-contained without relying on the original larger projects. After applying these criteria, we constructed 20 representative projects per programming language. The summary of dataset statistics is presented in Table 1, and detailed information on the dataset sources and statistics for each project can be found in Appendix A.
Pre-processing. Dataset pre-processing involves several key steps to ensure the projects are well-structured and suitable for testing. First, we double-check whether the selected projects have syntax errors, even though they are sourced from reliable codebases. Second, for projects extracted from a larger codebase, we modify them to be self-contained by reorganizing files, adjusting domain naming conventions, and/or modifying import paths to remove dependencies on external modules. Next, to enhance the accuracy of line coverage measurements, we consolidate statements that are split across multiple lines into a single line, ensuring that the metrics are more valid. Additionally, we maintain the integrity of the original code style as much as possible, preserving the diverse coding practices across different projects. This approach allows us to test how LLMs handle various code styles in a realistic environment.
@@ 3.2 Evaluation Metrics
We focus on three key aspects when evaluating the generated unit tests: compilation rate, correctness rate, and coverage rate. Compilation rate (ComR) measures the percentage of projects in which the generated test suites compile successfully, indicating how often LLMs produce unit test suites that can be executed without compilation errors. The compilation rate for all projects in $X$ is defined as $ComR=\frac{|X^{com}|}{|X|}$, where $X$ is the project set and $X^{com}\subset X$ denotes the subset of projects whose test suites compile successfully. Correctness rate (CR) calculates the percentage of unit tests that are correct out of all generated unit tests for each project, providing insight into the accuracy of the test generation process. On average, more than 95% of vanilla-generated unit tests compare expected and actual values, reinforcing the validity of CR as an evaluation metric. Detailed statistics see Appendix C. The correctness rate for the project $x$ is defined as $CR_{x}=\frac{|T_{x}^{cor}|}{|T_{x}|}$, where $T_{x}$ is the generated test suite and $T_{x}^{cor}\subset T_{x}$ denotes the correct unit test set for the project $x$. Coverage rate analyzes both line and branch coverage to understand how well the generated unit tests explore the code’s functionality. The coverage rate for the project $x$ is defined as $CR_{x}=\frac{covered(x)}{total(x)}$, where $covered(x)$ denotes the number of covered lines/branches in project $x$ and $total(x)$ the total number of lines/branches in project $x\in X$.
@@ 3.3 Unit Test Generation
Scenario 3: LLM Self-fixing. Inspired by our observation from manual fixing that different LLMs exhibit significantly different potentials after manual fixing, we seek to investigate how LLMs perform in self-fixing on our benchmark. We explore LLMs’ self-fixing abilities by incorporating conversation history and error messages as shown in Figure 6. We provide LLMs with the conversation history (including the system prompt, the user prompt for unit test generation requests, and LLM vanilla response), error messages obtained from the testing framework, and the user prompt for error fixing requests. When the open-source LLM’s input length is limited, we prioritize the information in the following order: system prompt, LLM’s initial response, error messages, error-fixing requests, and unit test generation requests. Less important information is truncated as needed. Additionally, we reserve at least 2,000 tokens for the open-source LLM’s self-fixing outputs. LLM self-fixing scenario helps us understand LLMs’ error-fixing ability and their potential to generate better unit tests when incorporating the self-fixing process. Note that during self-fixing, we do not constrain the target error types to just compilation or cascade errors.
@@ 5.1 Main Results
Table 2 shows the evaluation results of the vanilla unit tests. First, we observe that different LLMs have varying language-level expertise. For example, Claude-3.5-Sonnet performs the best in Java but falls behind GPT-o1 in JavaScript. Second, we can see from the results that LLMs have different metric-level expertise as well, validating the effectiveness of different evaluation metrics. For example, in Python, Claude-3.5-Sonnet performs the best on CR and ComR while falling behind GPT-o1 on LC and BC.
Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Many models fail to generate valid Java code, leading to low compilation rates and execution coverage. Among all the evaluated models, GPT-o1 performs the best in general, especially in JavaScript. CodeLlama and CodeGemma have the worst general performance. We also observe that some models tend to generate more unit tests. However, generating more unit tests does not necessarily lead to better coverage rates. For example, Gemini-1.5-Pro tends to generate the most unit tests but does not obtain the best coverage rate. Additionally, we observe that sometimes the open-source model can even outperform some closed-source models. For example, DeepSeek-Coder works better than GPT-3.5-Turbo on Python and JavaScript. Finally, we confirmed from such results that dependencies exist in metrics. On Java, models like CodeQwen1.5, CodeLlama, and CodeGemma fail to generate compilable unit tests, resulting in the lowest correctness rates and coverage rates.
@@ 5.4 Unique Contribution of Unit Tests
We also explore the unique contribution of the generated unit tests on Python. The unique contribution is defined as the total portion of coverage contributed by each generated unit test that does not overlap with the coverage of other unit tests. This is important for several reasons. First, some LLMs generate more unit tests than others, making it insufficient to rely solely on coverage rate as a metric; the unique contribution of each test should also be considered. Second, it is crucial for LLMs to generate fewer unit tests while still achieving a high coverage rate, as running a large number of tests can sometimes be resource- or time-intensive.